# Extended formulations for induced tree and path polytopes of chordal graphs

### v2

@@ 2.1 Induced paths of chordal graphs and Hilbert basis
Suppose that $\Psi^{u}$ is nonempty. $\mathbf{\Psi^{u}}$-Reduction is valid.
Proof. Given $\in\Psi^{u}, PP$ ends in $u$ and does not intersect $N(v)\setminus\{u\}$, therefore $P\cup\{v\}$ induces a path. Moreover, $\xi^{P}|_{H}=\xi^{P\cup\{v\}}|_{H}$. ∎
@@
Suppose that $\Psi_{K_{\Psi}}^{u}$ is non-empty. $\mathbf{\Psi_{K_{\Psi}}^{u}}$-Reduction is valid and strictly increments $\lambda(\Psi^{u}_{2})$, $\lambda(\Psi^{v}_{u})$.
Proof. By construction, $P\in\Psi_{K_{\Psi}}^{u}$ is decomposed into its two inclusion-wise maximal subsets $P_{1}$ and $P_{2}$ that induce paths ending in $u$. At most one of them intersects $N(v)\setminus\{u\}$, which implies that $P_{1}\cup\{v\}$ induces a path. Moreover, $\xi^{P}|_{H}=\xi^{P_{1}\cup\{v\}}|_{H}+\xi^{P_{2}}|_{H}.$ Therefore, $\mathbf{\Psi_{K_{\Psi}}^{u}}$-Reduction is valid and increments both $\lambda_{P_{1}\cup\{v\}}$ and $\lambda_{P_{2}}$. ∎
@@
Suppose that $\Psi_{\overline{K}_{\Psi}}^{u}$ contains an element $P$ such that a neighbor $u_{1}\in P$ of $u$ is not complete to $K_{\Psi}$, in particular $\Psi_{2}^{u}$ is non empty. Then $\mathbf{\Psi_{\overline{K}_{\Psi}}^{u}}$-Reduction is valid.
Proof. Let $P_{1}$ be the inclusion-wise maximal subset of $P$ inducing a path ending in $u$ and containing $u_{1}$, let $P^{\prime}$ be its other subset inducing a path ending in $u$. Let $P_{2}$ be an element of $\Psi^{u}_{2}$ such that the neighbor $u_{2}$ of $u$ in $P_{2}$ is non adjacent to $u_{1}$. Claim 2.2 implies that $P_{1}\cup P_{2}$ induces a path. Moreover, $\xi^{P}|_{H}+\xi^{P_{2}}|_{H}=\xi^{P_{1}\cup P_{2}}|_{H}+\xi^{P^{\prime}\cup\{v\}}|_{H}.$ Therefore $\mathbf{\Psi_{\overline{K}_{\Psi}}^{u}}$-Reduction is valid. ∎
@@
Suppose that $\Psi_{u}^{v}$ contains an element $P$ such that the neighbor $u_{1}$ of $u$, distinct from $v$, is not complete to a vertex $u_{2}$ of $K_{\Psi}$ contained in the path $P_{2}\in\Psi_{2}^{u}$. $\mathbf{\Psi_{u}^{v}}$-Reduction is valid and strictly increases $\lambda_{\{u,v\}}$.
Proof. Let $P_{1}=P\setminus\{v\}$, by Claim 2.2 $P_{1}\cup P_{2}$ induces a path and $\{u,v\}$ induces a path. Moreover, $\xi^{P}|_{H}+\xi^{P_{2}}|_{H}=\xi^{P_{1}}|_{H}+\xi^{\{u,v\}}|_{H}$. ∎
@@
Suppose that $\Psi_{\overline{K}_{\Psi}}^{u}$ is non empty. $\mathbf{\lambda_{\{u,v\}}}$-Reduction is valid and strictly decreases $\lambda_{\{u,v\}}$.
Proof. Given $P\in\Psi_{\overline{K}_{\Psi}}^{u}$, we consider its two inclusion-wise maximal subsets $P_{1}$ and $P_{2}$ inducing a path ending in $u$. $P_{1}\cup\{v\}$ induces a path. Moreover, $\xi^{P}|_{H}+\xi^{\{u,v\}}|_{H}=\xi^{P_{1}\cup\{v\}}|_{H}+\xi^{P_{2}}|_{H}$. ∎
The proof of the following theorem proceeds by induction over the number of vertices and makes use of the previous definitions and operations.
@@
The extended incidence vectors of induced paths of a chordal graph form a Hilbert basis of $\mathcal{C}_{P}(G)$.
Proof. The statement is trivial for $|V|=1$. Induction goes as follows: As $G$ is chordal, it contains a simplicial vertex $v$. The graph $G\setminus v$ is chordal and by hypothesis, the induced paths’ extended incidence vectors of $G\setminus v$ form a Hilbert basis of $\mathcal{C}_{P}(G\setminus v)$. Let $(x^{*},y^{*})$ belong to $\mathcal{C}_{P}(G)\cap\mathds{Z}^{V\cup E}$, its restriction $(x^{*},y^{*})|_{G\setminus v}$ belongs to $\mathcal{C}_{P}(G\setminus v)\cap\mathds{Z}^{V\cup E}$. From now on, $(\Psi,\lambda)$ denotes an arbitrary family of induced paths of $G$ with integer coefficients satisfying $(x’,y’)|_{G\setminus v}=(x^{*},y^{*})|_{G\setminus v}$, where (x,y)(x’,y’) is the point induced by $(\Psi,\lambda)$. All operations defined above transform $(\Psi,\lambda)$ into another such family and preserve: integrality of $\lambda$, the fact that the elements of $\Psi$ induce paths of $G$ and $(x’,y’)|_{H}=(x^{*},y^{*})|_{H}$. Note that the latter requirement implies that the operations not only preserve the value of $\lambda(\Psi)$ over elements in $G\setminus v$ but also on edges in $\delta(v)\setminus\{uv\}$.
The goal of the upcoming sequence of reductions is to achieve $(x^{\prime},y^{\prime})=(x^{*},y^{*})$. For this we consider the slackness;
$\Delta:=\sum_{uv\in\delta(v)}\Delta(u,v)\text{ where }\Delta(u,v)=y_{uv}^{*}-y_{uv}^{\prime}.$
The proof proceeds by fixing a neighbor $u$ of $v$ and modifying $(\Psi,\lambda)$ until the equality $\Delta(u,v)=0$ holds. Since reductions for a fixed $u$ do not affect $\Psi_{w}^{v}$ for $w\neq u$, and that elements of $\Psi^{u}_{w}$ are the only ones contributing to the value of $y_{vw}^{\prime}$, the procedure can be applied independently for each neighbor of $v$.
An induced path $P$ is removed from $\Psi$ if $\lambda_{P}$ is set to 0.
Lemmas 2.32.5 give conditions for some valid reductions strictly decreasing $\Delta(u,v)$ to be applied for respectively $\alpha=\min(\Delta(u,v),\lambda_{P})$, $\alpha=\min(\Delta(u,v),\lambda_{P})$, and $\alpha=\min(\lambda_{P},\lambda_{P_{2}},\Delta(u,v))$. Since the corresponding reductions strictly decrease either $\Delta(u,v)$, at some point, none of Lemma 2.32.6 apply for the fixed neighbor $u$. This implies that, either $\alpha=0$ and $\Delta(u,v)=0$ which ends the proof or the following holds:
  • $\Psi^{u}$, $\Psi^{u}_{I_{1}}$ both are empty
  • the neighbors of $u$ in the elements of $\Psi^{u}_{I_{0}}$ are complete to $K_{\Psi}$.
By construction, the only elements of $\Psi$ which contribute to the value of $y_{uv}^{\prime}$ belong to $\Psi_{u}^{v}$, therefore the following holds :
$y_{uv}^{\prime}=\lambda(\Psi_{u}^{v}).$ (8)
As $\Psi^{u}$ is empty, the following holds:
$x^{*}_{u}=x_{u}^{\prime}=\lambda(\Psi_{2}^{u})+\lambda(\Psi_{\overline{K}_{\Psi}}^{u})+\lambda(\Psi_{u}^{v})\\$ (9)
Suppose that $\Psi^{u}_{I_{0}}$ is empty. As $(x^{*},y^{*})$ belongs to $\mathcal{C}_{P}(G)$, by inequality (5) associated with $u$ and $K=N[v]\setminus\{u\}$ we get: $x_{u}^{*}\geq y^{*}(\delta(u,K))$ and the right hand side is larger than or equal to $\lambda(\Psi_{2}^{u})+y_{uv}^{*}$. Using equalities (8) and (9) we get : $y_{uv}^{\prime}\geq y_{uv}^{*}$ and hence $\Delta(u,v)=0$.
Assume henceforth that $\Psi_{\overline{K}_{\Psi}}^{u}$ is non-empty.
Suppose $\Psi_{2}^{u}$ to be empty, Inequality (5) associated with $u$ and $K=\emptyset$ applied to $(x^{*},y^{*})$ yields: $2x_{u}\geq y^{*}(\delta(u))$. Note that the right hand side is larger than or equal to $2\lambda(\Psi_{I}^{u})+\lambda(\Psi_{u}^{v})+y_{uv}^{*}$ as each $\lambda_{P}$ for $P\in\Psi_{I}^{u}\cup\Psi^{v}_{u}$ contributes two times in $y^{*}(\delta(u))$ and $\lambda(\Psi^{v}_{u})=y_{uv}^{*}$. Equalities (9) and (8) yield $y_{uv}^{\prime}\geq y_{uv}^{*}$ and $\Delta(u,v)=0$.
We may now assume that $\Psi_{2}^{u}$ is nonempty.
Therefore, $K$ is non-empty and the neighbors of $u$ in the elements of $\Psi_{\overline{K}_{\Psi}}^{u}$ are complete to $K_{\Psi}$.
Suppose that the conditions of Lemma 2.6 hold, then applying $\Psi_{u}^{v}$-Reduction with $\alpha=\min(\lambda_{P},\lambda_{P_{2}})$ yields an element of $\Psi_{K_{\Psi}}^{u}$, therefore, the conditions of Lemma 2.4 hold and $\Psi_{K_{\Psi}}^{u}$-Reduction applies. This sequence of operations strictly decreases $\Delta(u,v)$.
Suppose that the conditions of Lemma 2.7 hold, applying $\lambda_{u}^{v}$-Reduction with $\alpha=\min(\lambda_{P},\lambda_{u,v}$) yields an element of $\Psi_{u}^{v}$, and one of $\Psi^{u}$. Therefore, the conditions of Lemma 2.3 hold, and $\Psi^{u}$-Reduction applies. This sequence of operations also decreases $\Delta(u,v)$.
As $\Psi^{u}$-Reduction, $\Psi_{K_{\Psi}}^{u}$-Reduction, $\Psi_{\overline{K}_{\Psi}}^{u}$-Reduction, and the two latter sequences of reductions strictly decrease $\Delta(u,v)$, we may assume that none of Lemmas 2.32.7 apply or that $\Delta(u,v)=0$. Therefore, the neighbors of $u$ in the elements of $\Psi^{v}_{u}$ are complete to $K_{\Psi}$ and $\lambda_{\{u,v\}}=0$.
As $(x^{*},y^{*})$ is a point of $\mathcal{C}_{P}(G)$, it satisfies the constraint (5) associated with $K_{\Psi}$ and $u$ that is (where $C(S)$ denotes the set of vertices complete to $S$):
$2x_{u}^{\prime}=2x_{u}^{*}\geq 2y^{*}(\delta(u,K_{\Psi}))+y^{*}(\delta(u,C(K_{\Psi}\cup\{u\})))$ (10)
@@
By construction, each element $P$ of $\Psi_{u}^{v}$ contains both $v$ and another neighbor of $u$ and both are complete to $K_{\Psi}$. This implies that $\lambda_{P}$ contributes two times in $y^{*}[\delta(u,C(K_{\Psi}\cup\{u\}))]$. By construction, $y^{*}(\delta(u,C(K_{\Psi}\cup\{u\})))=y^{*}_{uv}+y^{*}[\delta(u,C(K_{\Psi}\cup\{u\})\setminus\{uv\})]=y^{*}_{uv}+\lambda(\Psi_{u}^{v})+2\lambda(\Psi_{\overline{K}_{\Psi}}^{u})$. Using equalities (8) and (9) we get:
$0\geq y^{*}_{uv}-\lambda(\Psi_{u}^{v})$
@@
which finally shows that $y_{uv}^{\prime}\geq y_{uv}^{*}$. Repeat this procedure until $y_{uv}^{*}=y_{uv}^{\prime}\ \forall uv\in\delta(v)$. Setting $\lambda_{\{v\}}=(x^{*}_{v}-y^{*}(\delta(v)))$ for the trivial induced path $\{v\}$ yields the desired integer combination of $(x^{*},y^{*})$ using induced paths’ extended incidence vectors.
@@
As induced trees of $C_{4}$ coincide with its induced path, Theorem 2.8 is not a characterization of the graph for which the induced path extended incidence vectors have IDP.
2.2 Extended formulation for the induced path polytope

### v3

@@ 2.1 Induced paths of chordal graphs and Hilbert basis
Given $u\in N(v)$, either $\Delta_{uv}(\Psi,\lambda)=0$ or $\Psi^{u}=\emptyset$.
Proof. By contradiction let $P$ be an element of $\Psi^{u}$ and $\alpha=\min\Bigl(\Delta_{uv}(\lambda,\Psi),\lambda_{P}\Bigr)$. Decreasing $\lambda_{P}$ and increasing $\lambda_{P\cup\{v\}}$ by $\alpha$ yields a new integer conic combination $(\lambda^{\prime},\Psi^{\prime})$, of induced paths’ extended incidence vectors. Moreover, as $\xi^{P}|_{G\setminus uv}=\xi^{P\cup v}|_{G\setminus uv}$, we have that $\Bigl(\sum_{P\in\Psi^{\prime}}\lambda_{P}^{\prime}\xi^{P}\Bigr)\Bigm\lvert_{G\setminus uv}=(x^{\prime},y^{\prime})|_{G\setminus uv}$. Therefore $(x^{\prime},y^{\prime})$ satisfies Conditions (i)-(ii), $\Delta_{e}(\lambda^{\prime},\Psi^{\prime})=\Delta_{e}(\lambda,\Psi)\ \forall e\in\delta(v)\setminus uv$ and $\Delta(\lambda^{\prime},\Psi^{\prime})=\Delta(\lambda,\Psi)-\alpha$. By Condition (iii) applied to $(\lambda,\Psi)$ we obtain $\alpha=0$ and as $\lambda_{P}>0$ we finally obtain $\Delta_{uv}(\Psi,\lambda)=0$. ∎
By Lemma 2.5, if $\Delta_{uv}(\Psi,\lambda)$ is nonzero, we have $\Psi^{u}=\emptyset$. Therefore, the following inequality holds and will be used in the upcoming results:
$x^{*}_{u}=x_{u}^{\prime}=\lambda(\Psi_{2}^{u})+\lambda(\Psi_{I_{2}}^{u})+\lambda(\Psi_{I_{1}}^{u})+\lambda(\Psi_{u}^{v}).\\$ (9)
@@
Given $u\in N(v)$, either $\Delta_{uv}(\lambda,\Psi)=0$ or $\Psi^{u}_{I_{1}}$ is nonempty.
Proof. Suppose that $\Delta_{uv}(\lambda,\Psi)\neq 0$ and that $\Psi^{u}_{I_{1}}=\emptyset$. As $(x^{*},y^{*})$ belongs to $\mathcal{C}_{P}(G)$, by inequality (5) associated with $u$ and $K=N[v]\setminus\{u\}$ we get: $x_{u}^{*}\geq y^{*}(\delta(u,K))$ and the right-hand side is equal to $\lambda(\Psi_{2}^{u})+\lambda(\Psi_{I_{2}}^{u})+y_{uv}^{*}$ by definition of $(\lambda,\Psi)$. Using equalities (8) and (9) we get : $y_{uv}^{\prime}\geq y_{uv}^{*}$ and hence $\Delta_{uv}(\lambda,\Psi)=0$, a contradiction. ∎
@@
Given $u\in N(v)$, the neighbors of $u$ in the elements of $\Psi_{I_{1}}^{u}$ are complete to $K_{2}^{u}\cup\{u\}$.
Proof. By contradiction, suppose that a neighbor $u_{1}$ of $u$ in some element $P\in\Psi^{u}_{I_{1}}$ is not adjacent to a vertex $u_{2}\in K_{2}^{u}$ contained in some induced path $P_{2}\in\Psi^{u}_{2}$. Let $\alpha=\min\Bigl(\Delta_{uv}(\lambda,\Psi),\lambda_{P},\lambda_{P_{2}}\Bigr)$. Let us consider the two inclusion wise maximal induced path contained in $P$ ending in $u$, say $_{1}, PP^{\prime}$, and let $P_{1}$ contain $u_{1}$. By Claim 2.2, $P_{1}\cup P_{2}$ induces a path of $G$. Decreasing $\lambda_{P}$, $\lambda_{P_{2}}$ and increasing $\lambda_{P_{1}\cup P_{2}}$, $\lambda_{P^{\prime}\cup\{v\}}$ yields a nonnegative integer combination $(\lambda^{\prime},\Psi^{\prime})$ of induced paths’ extended incidence vectors. Moreover, the following equality holds $\Bigl(\xi^{P}+\xi^{P_{2}}\Bigr)\Bigm\lvert_{G\setminus uv}=\Bigl(\xi^{P_{1}\cup P_{2}}+\lambda_{P^{\prime}\cup\{v\}}\Bigr)\Bigm\lvert_{G\setminus uv}$. Therefore, $(\lambda^{\prime},\Psi^{\prime})$ satisfies Conditions (i)-(ii), $\Delta_{e}(\lambda^{\prime},\Psi^{\prime})=\Delta_{e}(\lambda,\Psi)\ \forall e\in\delta(v)\setminus\{uv\}$ and $\Delta_{uv}(\lambda^{\prime},\Psi^{\prime})=\Delta_{uv}(\lambda,\Psi)-\alpha$. We finally get $\alpha=0$ and hence $\Delta_{uv}(\lambda,\Psi)=0$. ∎
@@
Given $u\in N(v)$ either $\Delta_{uv}(\lambda,\Psi)=0$ or $\lambda_{u,v}=0$.
Proof. Suppose by contradiction that both $\lambda_{u,v}$ and $\Delta_{uv}(\lambda,\Psi)$ are nonzero. By Lemma 2.6 $\Psi_{I_{1}}^{u}$ is nonempty, let $P$ be an element of $\Psi_{I_{1}}^{u}$. $P$ contains the two distinct inclusion-wise maximal subsets $P_{1}$ and $P_{2}$ both inducing paths ending in $u$. Let $\alpha=\min(\Delta_{uv}(\lambda,\Psi),\lambda_{P},\lambda_{\{u,v\}})$. Note that $P_{1}\cup\{v\}$ induces a path. Decreasing $\lambda_{u,v}$, $\lambda_{P}$ and increasing $\lambda_{P_{2}}$, $\lambda_{P_{1}\cup\{v\}}$ by $\alpha$ yields a nonnegative integer combination $(\lambda^{\prime},\Psi^{\prime})$. As $\xi^{\{u,v\}}+\xi^{P}=\xi^{\{u,v\}\cup P_{1}}+\xi^{P_{2}}$, $(\lambda^{\prime},\Psi^{\prime})$ satisfies Conditions (i)-(ii) and $\Delta_{uv}(\lambda,\Psi)=\Delta_{uv}(\lambda^{\prime},\Psi^{\prime})$. However, $(\Psi^{\prime})^{u}$ is nonempty as it contains $P_{2}$, by Lemma 2.5, $\Delta_{uv}(\lambda,\Psi)=0$, a contradiction. ∎
@@
Given $u\in N(v)$, either $\Delta_{uv}(\lambda,\Psi)=0$ or $K_{2}^{u}$ is nonempty.
Proof. Suppose by contradiction that $\Psi_{2}^{u}$ is empty and that $\Delta_{uv}(\lambda,\Psi)$ is strictly positive. By Lemma 2.5, Equality (9) holds. Inequality (5) associated with $u$ and the empty clique applied to $(x^{*},y^{*})$ yields: $2x_{u}^{*}\geq y^{*}(\delta(u))$. By Lemma 2.8, $\lambda_{u,v}=0$. Therefore, each $\lambda_{P}$ for $P\in\Psi_{I_{1}}^{u}\cup\Psi_{I_{2}}\cup\Psi^{v}_{u}$ contributes two times in $y^{*}(\delta(u))$, hence the following holds:
$2x_{u}^{*}\geq 2\Bigl(\lambda(\Psi_{I_{1}}^{u})+\lambda(\Psi_{I_{2}}^{u})\Bigr)+\lambda(\Psi_{u}^{v})+y_{uv}^{*}.$
Equalities (8)-(9) yields $y_{uv}^{\prime}\geq y_{uv}^{*}$ implying that $\Delta_{uv}(\lambda,\Psi)=0$, a contradiction. ∎
Lemma 2.10.
Given $u\in N(v)$, either $\Delta_{uv}(\lambda,\Psi)=0$ or the neighbors of $u$ in the elements of $\Psi_{u}^{v}$ are complete to $K_{2}^{u}\cup\{u\}$.
Proof. Suppose that $\Delta_{uv}(\lambda,\Psi)$ is nonzero. By Lemma 2.8, the elements of $\Psi_{u}^{v}$ have at least 3 vertices. Let $P=(v,u,u_{1},\dots)$ be an element of $\Psi^{v}_{u}$. By construction, $v$ is complete to $K_{2}^{u}\cup\{u\}$. Suppose that $u_{1}$ is not complete to $K_{2}^{u}\cup\{u\}$ and, without loss of generality, not adjacent to the vertex $u_{2}\in P_{2}\cap K_{2}^{u}$ for some induced path $P_{2}\in\Psi_{2}^{u}$. Let $P_{1}$ be the inclusion-wise maximal induced path included in $P$ ending in $u$ and containing $u_{1}$. By Claim 2.2, $P_{1}\cup P_{2}$ induces a path. Let $\alpha=\min(\lambda_{P},\lambda_{P_{2}})$, decreasing $\lambda_{P}$, $\lambda_{P_{2}}$ and increasing $\lambda_{P_{1}\cup P_{2}}$, $\lambda_{\{u,v\}}$ by $\alpha$ yields a nonnegative integer combination $(\lambda^{\prime},\Psi^{\prime})$ of induced paths’ extended incidence vectors. As $\xi^{P}+\xi^{P_{2}}=\xi^{P_{1}\cup P_{2}}+\xi^{\{u,v\}}$, $(\lambda^{\prime},\Psi^{\prime})$ satisfies Conditions (i)-(iii), and $\Delta_{uv}(\lambda,\Psi)=\Delta_{uv}(\lambda^{\prime},\Psi^{\prime})$. Moreover, we have that $\lambda^{\prime}_{\{u,v\}}$ is nonzero, therefore by Lemma 2.8, $\Delta_{uv}(\lambda,\Psi)=0$. ∎
@@
The extended incidence vectors of induced paths of a chordal graph $G$ form a Hilbert basis of $\mathcal{C}_{P}(G)$.
Proof. Similarly to the proof of Theorem 1.1, the first step is to show that $\Delta_{e}(\lambda,\Psi)$ is equal to 0 for every $e\in\delta(v)$ for $(\lambda,\Psi)$ satisfying Conditions (i)-(iii). This is done through the use of an inequality (5) associated with $u$. Therefore, we need to exhibit the right clique of $N(u)$. We iteratively construct this clique that we denote by $K_{\Psi}$, and note $\Psi_{K_{\Psi}}$ the subset of $\Psi^{u}_{2}\cup\Psi_{I_{2}}^{u}$ whose elements contain a vertex of $K_{\Psi}$. Note that $\Psi_{K_{\Psi}}$ is supposed to be changing along to $K_{\Psi}$.
Start with $K_{\Psi}=K_{2}^{u}$. If there exists an element $P$ of $\Psi_{I_{2}}^{u}\setminus\Psi_{K_{\Psi}}$ such that the vertex $w\in\Bigl(P\cap N(u)\Bigr)\setminus N(v)$ is not complete to $K_{\Psi}$, add $P$ to $\Psi_{K_{\Psi}}$ and add the other neighbor of $u$ in $P$ to $K_{\Psi}$. This procedure is applied until the neighbors of $u$ in the elements of $\Psi_{I_{2}}^{u}\setminus\Psi_{K_{\Psi}}$ are complete to $K_{\Psi}\cup\{u\}$.
By Lemmas 2.7-2.10, the neighbors of $u$ in the elements of $\Psi_{I_{1}}^{u}\cup\Psi_{u}^{v}$ are complete to $K_{2}^{u}\cup\{u\}$. Moreover, by the previous construction, the elements of $\Psi^{u}_{I_{2}}\setminus\Psi_{K_{\Psi}}$ satisfy this property for the whole clique $K_{\Psi}\cup\{u\}$.
It remains to show that the neighbors of $u$ in the elements of $\Psi_{I_{1}}^{u}\cup\Psi_{u}^{v}$ are complete to $K_{\Psi}\setminus K_{2}^{u}$. We do so by reorganizing the elements of $\Psi_{K_{\Psi}}$.
Each vertex of $K_{\Psi}\setminus K_{2}^{u}$ is contained in an element of $\Psi_{I_{2}}^{u}\cap\Psi_{K_{\Psi}}$. Let $P$ be an element of $\Psi_{I_{1}}^{u}\cup\Psi^{v}_{u}$ and let $P^{\prime}$ be an element of $\Psi_{K_{\Psi}}\setminus\Psi_{2}^{u}$. As $P^{\prime}$ has been added to $\Psi_{K_{\Psi}}$ by the previous construction, there exists a sequence of induced path, say $(P_{1},\dots,P_{\ell})$ such that $P_{1}=P^{\prime}$, $P_{\ell}\in\Psi_{2}^{u}$ and all the other elements belong to $\Psi_{I_{2}}^{u}$, such that the vertex $u_{i}\in(N(u)\cap P_{i})\setminus K_{\Psi}$ is nonadjacent to the vertex $v_{i+1}\in K_{\Psi}\cap P_{i+1}\ \forall i\in\{1,\dots,\ell-1\}$. For each $P_{i},\ i\in\{1,\dots,\ell-1\}$, we consider the two maximal induced path $U_{i},V_{i}$ included in $P_{i}$ and ending in $u$, with $U_{i}$ containing $u_{i}$ and $V_{i}$ containing $v_{i}$, we set $V_{\ell}=P_{\ell}$. Claim 2.2 implies that $U_{i}\cup V_{i+1}\ \forall i\in\{1,\dots,\ell-1\}$ induces a path, see Figure 4 for an illustration where the dashed line represents a nonedge. Therefore, for $\alpha=min\{\lambda_{P_{i}}:i\in\{1,\dots,\ell\}\}$ decreasing $\lambda_{P_{i}}\ \forall i\in\{1,\dots,\ell\}$, increasing $\lambda_{V_{i+1}\cup U_{i}}\forall i\in\{1,\dots,\ell\-1\}$ and $\lambda_{V_{1}},$ by $\alpha$ yields a nonnegative combination of induced paths’ extended incidence vectors $(\lambda^{\prime},\Psi^{\prime})$. Moreover $\sum_{i=1}^{\ell}\xi^{P_{i}}=\xi^{V_{1}}+\sum_{i=1}^{\ell-1}\xi^{U_{i}\cup V_{i+1}}$, implying that, $(\lambda^{\prime},\Psi^{\prime})$ satisfies Conditions (i)-(iii) and $(\Psi^{\prime})^{u}_{2}$ contains $V_{1}$. By Lemmas 2.7 and 2.10, $P$ contains two vertices adjacent to the only vertex in $V_{1}\cap K_{\Psi}$. Note that this observation can be made for each element in $\Psi_{K_{\Psi}}\setminus\Psi_{2}^{u}$ and each element in $\Psi^{u}_{I_{1}}\cup\Psi^{v}_{u}$. This implies that the neighbors of $u$ in the elements of $\Psi^{u}_{I_{1}}\cup\Psi^{v}_{u}$ have two vertices complete to $K_{\Psi}\cup\{u\}$.
As $(x^{*},y^{*})$ is a point of $\mathcal{C}_{P}(G)$, it satisfies the constraint (5) associated with $K_{\Psi}$ and $u$ that is:
$2x_{u}^{\prime}=2x_{u}^{*}\geq y^{*}\Biggl(\delta\Bigl(u,C\left(K_{\Psi}\cup\{u\}\right)\Bigl)\Biggr)+2y^{*}\Biggl(\delta\Bigl(u,K_{\Psi} \Bigr)\Biggr)$ (10)
@@
By the previous paragraph, each element $P$ of $\Psi_{u}^{v}\cup\Psi_{I_{1}}^{u}\cup\Psi_{I_{2}}^{u}\setminus\Psi_{K_{\Psi}}$ contains $u$ and its neighbors in $P$ are complete to $K_{\Psi}$. This implies that $\lambda_{P}$ contributes two times in
$y^{*}\Biggl(\delta\Bigl(u,C(K_{\Psi}\cup\{u\})\Bigr)\Biggr).$
Therefore, the following holds:
$\begin{array}[]{rl}y^{*}\Biggl(\delta\Bigl(u,C(K_{\Psi}\cup\{u\})\Bigr)\Biggr)&=y^{*}_{uv}+y^{*}\Biggl(\delta\Bigl(u,C(K_{\Psi}\cup\{u\})\Bigr)\setminus\{uv\}\Biggr)\\ &=y^{*}_{uv}+\lambda(\Psi_{u}^{v})+2\lambda(\Psi_{I_{1}}^{u})+2\lambda(\Psi_{I_{2}}^{u}\setminus\Psi_{K_{\Psi}}).\end{array}$
Moreover, each element of $\Psi_{K_{\Psi}}$ intersects $K_{\Psi}$ and contains $u$, implying that it contributes one time in $y^{*}\Biggl(\delta\Bigl(u,K_{\Psi} \Bigr)\Biggr)$. Using equality (9) and inequality (10) we get:
$2\Bigl(\lambda(\Psi_{2}^{u})+\lambda(\Psi_{I_{2}}^{u})+\lambda(\Psi_{I_{1}}^{u})+\lambda(\Psi_{u}^{v})\Bigr)\geq y^{*}_{uv}+\lambda(\Psi_{u}^{v})+2\lambda(\Psi_{I_{1}}^{u})+2\lambda(\Psi_{I_{2}}^{u}\setminus\Psi_{K_{\Psi}})+2\Bigl(\lambda(\Psi_{K_{\Psi}})\Bigr)$
@@
Moreover, $\Psi_{I_{2}}^{u}=(\Psi_{I_{2}}^{u}\cap\Psi_{K_{\Psi}})\cup(\Psi_{I_{2}}^{u}\setminus\Psi_{K_{\Psi}})$ and $\Psi_{I_{2}}^{u}\cap\Psi_{K_{\Psi}}=\Psi_{K_{\Psi}}\setminus\Psi_{2}^{u}$. Therefore, $y_{uv}^{\prime}\geq y_{uv}^{*}$ by equality (8). This reasoning can be made for each $uv\in\Delta(v)$. Therefore, $\Delta(\lambda,\Psi)=0$. Setting $\lambda_{\{v\}}=(x^{*}_{v}-y^{*}(\delta(v)))$ for the trivial induced path $\{v\}$ yields the desired nonnegative integer combination of $(x^{*},y^{*})$ with induced paths’ extended incidence vectors. ∎
[…]
Figure 4: Structure of the paths exhibited in Theorem 2.11
@@
Remember that $C_{4}$ is a non-chordal graph for which the extended incidence vectors of induced trees form a Hilbert basis. Therefore, as induced trees of $C_{4}$ coincide with its induced path, Theorem 2.11 is not a characterization of the graph for which the induced path extended incidence vectors form a Hilbert basis. This result also raises the question of a characterization of the graphs for which induced paths’ extended incidence vectors form Hilbert bases.
2.2 Induced paths of chordal graphs and integral systems